OpenStudy (boom):

9^10n-12=1/6561 I'm stuck

OpenStudy (callisto):

9^(10n-12)=1/6561 9^(10n-12) = 9^(-4) 10n - 12 = -4 n = ... ?

OpenStudy (boom):

thanks